OpenStudy (anonymous):

2x^3-x^2+2x=0 Solve.

OpenStudy (anonymous):

To solve, you can factor. What is the term that can be taken out of all three of these?

OpenStudy (anonymous):

Oops, meant to write: 2x^2-x^2+2x-3=0. Just realized this can be done by group factoring right?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Im getting x\[x ^{2}(2x-1)(2x-3)=0. Somehow I can't get the two binomials in the parenthesis to be the same

OpenStudy (anonymous):

\[x ^{2}(2x-1)(2x-3)=0\]

OpenStudy (anonymous):

Same here.. o_o Hold on just a sec, im going to try something else

OpenStudy (anonymous):

Could be a teacher's mistake I guess

OpenStudy (anonymous):

Oh is it 2x^2, or 2x^3?

OpenStudy (anonymous):

2x^3

OpenStudy (anonymous):

This is what i got from Wolfram Alpha: http://www.wolframalpha.com/input/?i=2x%5E3-x%5E2%2B2x-3%3D0